2a² = 5a – 3 and 2b² = 5b – 3. What is the number value of a² + b² ?

Grade:10

3 Answers

Arun
4 years ago
 
Dear student
 
x^2 – 5 x + 3 = 0 has roots a and b.
 
hence a^2 + b^2 = (a + b)^2 – 2 ab = 25 – 6 = 19
Hope it helps
 
Regards
Arun
Aditya Gupta
4 years ago
hello sayantan, aruns ans is wrong.
note that a and b are roots of eqn 2x^2 – 5x+3= 0
so a^2 + b^2 = (a + b)^2 – 2 ab = 25/4 – 2*3/2
13/4
kindly approve :)
Arun
4 years ago
Dear student
 
Sorry, I forgot to take 2 in consider 
 
it is 2x^2 – 5x + 3 = 0
 
now a^2 + b^2 = (a + b)^2 – 2ab = 25/4 – 3 = 13/4
 
Thanks Aditya for making me correct